OpenStudy (anonymous):

Point A is located at (-2, 2), and D is located at (4, -2). Find the coordinates of the point that lies halfway between A and D. (-1, 1) (0, 1) (1, 0) (1, 1)

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

use mid point formula do u have it ?

ganeshie8 (ganeshie8):

mid point = \(\huge \left( \frac{x_1 + x_2 }{2} , \frac{y_1 + y_2}{2} \right)\)

ganeshie8 (ganeshie8):

(-2, 2), and D is located at (4, -2). x1 y1 x2 y2

ganeshie8 (ganeshie8):

simply plug them

OpenStudy (anonymous):

ok hold on

ganeshie8 (ganeshie8):

take ur time

OpenStudy (anonymous):

i got (1,0) which is c

ganeshie8 (ganeshie8):

Good work !